### v1

# Artificial Precision Timing Array: bridging the decihertz gravitational-wave sensitivity gap with clock satellites

Abstract
Gravitational-wave astronomy has developed enormously over the last decade with the first detections across different frequency bands, but has yet to access 0.1100.1100.1–100.1 – 10 Hz gravitational waves. Gravitational waves in this band are emitted by some of the most enigmatic sources, including intermediate-mass binary black hole mergers, early inspiralling compact binaries, and possibly cosmic inflation. To tap this exciting band, we propose the construction of a detector based on pulsar timing principles, the Artificial Precision Timing Array (APTA). We envision APTA as a solar system array of artificial “pulsars”—precision-clock-carrying satellites that emit pulsing electromagnetic signals towards Earth or other centrum. In this fundamental study, we estimate the clock precision needed for APTA to successfully detect gravitational waves. Our results suggest that a clock relative uncertainty of $10^{-17}$, which is currently attainable, would be sufficient for APTA to surpass LISA’s sensitivity in the decihertz band and observe $10^{3}–10^{4}$ $\mathrm{M}_{\odot}$ black hole mergers. Future atomic clock technology realistically expected in the next decade would enable the detection of an increasingly diverse set of astrophysical sources, including stellar-mass compact binaries that merge in the LIGO-Virgo-KAGRA band, extreme-mass-ratio inspirals, and Type Ia supernovae. This work opens up a new area of research into designing and constructing artificial gravitational-wave detectors relying on the successful principles of pulsar timing.
I Introduction
Since the first successful detection of gravitational waves (GWs) by LIGO in 2015, more than 100 compact-binary mergers have been seen by ground-based detectors Aasi et al. (2015); Abbott et al. (2016, 2021); Accadia et al. (2012). Moreover, the global network of pulsar timing arrays (PTAs) has provided strong evidence of the stochastic GW background Agazie et al. (2023a); Antoniadis et al. (2023); Reardon et al. (2023); Zic et al. (2023); Xu et al. (2023). With the fourth LIGO-Virgo-KAGRA observing run in progress and the continued PTA observations Manchester et al. (2013); Desvignes et al. (2016); Antoniadis et al. (2022); Agazie et al. (2023a), the era of gravitational-wave astronomy is surely underway.
Current techniques access a limited window of the GW frequency spectrum. The ongoing PTA observation campaign is sensitive in the $10^{-9}–10^{-6}$ Hz band Maiorano et al. (2021), with its upper-frequency limit set by its fortnightly access to radio telescopes. Meanwhile, LIGO-Virgo-KAGRA interferometers are sensitive in the $\mathrm{10\ Hz–10\ kHz}$ band. In the future, space-based interferometers such as Laser Interferometer Space Antenna (LISA) and TianQin may be sensitive to 0.11000.11000.1–1000.1 – 100 mHz GWs, while next-generation Earth-based missions such as Cosmic Explorer and Einstein Telescope will push the lower limit of ground-based sensitivity to a few Hz Amaro-Seoane et al. (2017); Huerta and Gair (2010). Seismic and gravity gradient noises, however, severely limit ground-based detector sensitivity below $\sim$ 1 Hz Abbott et al. (2009); Punturo et al. (2010); Abbott et al. (2016). Consequently, the 0.110.110.1–10.1 – 1 Hz remains inaccessible, although space-based interferometer concepts, such as the DECi-hertz Interferometer Gravitational wave Observatory (DECIGO) Kawamura et al. (2006) and the Big Bang Observer (BBO) Phinney et al. (2004); Harry et al. (2006), have been proposed.
Observational access to the 0.1–10 Hz GW band is imperative for the progress of many areas of astrophysics and cosmology. For example, $10^{3}–10^{4}$ $\mathrm{M}_{\odot}$ binary black holes have yet to be unambiguously observed. The observation of these sources can advance our understanding of supermassive black hole and galaxy formation, stellar clusters, and active galactic nuclei (AGN) Miller and Colbert (2004); Milosavljević et al. (2009); McKernan et al. (2012); Ferrara et al. (2014); Yang et al. (2019). While LISA may detect the early inspirals of intermediate-mass binaries Jani et al. (2020), a 0.1–10 Hz detector accesses the wealth of additional information only available during a merger. Other promising phenomena that may feature in this sensitivity gap include primordial GWs from inflation Corbin and Cornish (2006); Smith et al. (2006); Kuroyanagi et al. (2009); Guzzetti et al. (2016), the evolution of stellar-mass binaries and merging triple black hole systems Arca Sedda et al. (2021); O’Leary et al. (2009); Samsing (2018); Gondán et al. (2018); Antonini et al. (2017); Samsing and D’Orazio (2018); Codazzo et al. (2023), supernovae Ott (2009); Seitenzahl et al. (2015), and parabolic encounters Kocsis et al. (2006); Codazzo et al. (2023), among others further discussed in Section II.
To reach this unexplored GW band, we propose a new detector concept: the Artificial Precision Timing Array (APTA), a network of satellites distributed around the Earth carrying high-precision clocks to act as artificial “pulsars”. Having satellite signals aimed at a small station (or stations) dedicated to their reception, APTA can avoid the cadence constraint of regular PTAs and achieve high sensitivity at much higher frequencies. Then, if the satellite clocks are timed at least every $\sim$ 0.1 s, APTA can reach high sensitivity to 0.1–10 Hz GWs for observations lasting 10 s or more.
The idea of using high-precision clocks in space to detect GWs has been proposed and applied before. The global positioning system (GPS) was used to constrain the strain amplitude of the GW background in the 0.01–1 Hz band to $h_{c}<4.8\times 10^{-12}(1~{}{\rm Hz}/f)$ Aoyama et al. (2014), in an application of the Doppler tracking method Estabrook and Wahlquist (1975); Armstrong (2006); Vutha (2015). Another study proposed the use of two satellites in heliocentric orbits equipped with optical lattice atomic clocks to detect GWs, using a technique similar to Doppler tracking Kolkowitz et al. (2016). Another proposed detector, the Space Atomic Gravity Explorer (SAGE), consists of two satellites carrying ultracold atomic strontium systems to either perform atom interferometry with the strontium or Doppler tracking utilizing Sr optical atomic clocks Tino et al. (2019). Nevertheless, the idea of equipping artificial “pulsars” in the solar system with high-precision clocks to detect GWs with pulsar-timing-type methods is unprecedented.
In this paper, we describe the groundwork for a solar system APTA targeting the 0.1–10 Hz GW frequency band. In Section II, we list and discuss several interesting astrophysical targets observable in this band. In Section III, we summarize the basics of PTA sensitivity Moore et al. (2015a). In Section IV, we present the concept of APTA in detail, including its sensitivity for different configurations and potential performance. In Section V, we briefly list and discuss error-introducing factors to be incorporated into a more realistic future model of APTA. We conclude in Section VI.
II Gravitational Wave Astrophysics with APTA
The GW frequency regime between LISA and Earth-based detectors probes a diverse set of astrophysical and cosmological sources Arca Sedda et al. (2020). Notably, $10^{3}–10^{4}$ $\mathrm{M}_{\odot}$ black hole binaries, which merge before entering the LIGO-Virgo-KAGRA band, remain rich for study. The nature of these sources provides unparalleled insight into the formation of supermassive black holes and galaxies as well as AGN and stellar cluster dynamics Miller and Colbert (2004); Milosavljević et al. (2009); McKernan et al. (2012); Ferrara et al. (2014).
Inspiraling stellar-mass compact binaries also emit GWs in this band before merging in the ground-accessible band. Low-frequency GW observations of these sources directly probe their formation and evolution Nishizawa et al. (2016, 2017); Arca Sedda et al. (2021); O’Leary et al. (2009); Samsing (2018); Gondán et al. (2018); Antonini et al. (2017). Long-duration tracking may probe BBH eccentricity (Arca Sedda et al., 2020; Liu et al., 2022) as well as improve effective spin and localization measurements. Similarly the formation of merging triple black hole systems may also be probed in this band Samsing and D’Orazio (2018); Codazzo et al. (2023). Excitingly, early detection of an inspiraling compact binary could trigger multimessenger precursor counterpart searches Tsang et al. (2012); Sullivan et al. (2023) as well as give advanced warning of short gamma-ray bursts (Narayan et al., 1992; Abbott et al., 2017a, b) and kilonovae (see Ref. Gottlieb et al. (2023) and references therein).
The characteristic GW strain of an astrophysical binary well before merger may be estimated as Jaffe and Backer (2003); Finn and Thorne (2000a)
$h_{c}\approx 4.1\times 10^{-20}\left(\frac{M}{10^{3}\text{ M}{{}_{\odot}}}\right)^{\frac{5}{6}}\left(\frac{f_{GW}}{1\text{ Hz}}\right)^{-\frac{1}{6}}\left(\frac{D}{1\text{ Gpc}}\right)^{-1}\frac{q^{\frac{1}{2}}}{(1+q)},$ (1)
where $M$ is the total mass, $f_{GW}$ is the gravitational wave frequency, $D$ is the luminosity distance to the source, and $q\leq 1$ is the mass ratio. The merger frequency for a BBH may be estimated as
$f_{GW}^{merger}\approx\frac{1}{\pi}\frac{c^{3}}{2^{\frac{3}{2}}GM}\approx 1.3\times 10^{4}\text{ Hz}\left(\frac{M}{1\text{ M}_{\odot}}\right)^{-1}.$ (2)
Heavy stellar-mass BBHs (total mass $\sim 100$ $\mathrm{M}_{\odot}$) at distances of $\sim 1$ Gpc will have strains in the range $h\sim 10^{-21}–10^{-20}$ in dHz band during early inspiral. The late inspirals and mergers of $10^{3}–10^{4}$ $\mathrm{M}_{\odot}$ BBHs should have even larger dHz GW strains. For instance, an intermediate-mass binary black hole merger with $D\sim 1$ Gpc, and total mass $\sim 1\times 10^{4}$ ${}_{\odot}$, which will merge at $f_{GW}\approx 1.3$ Hz, will have strains exceeding $h\gtrsim 10^{-19}$. Since the GW frequency is half the orbital frequency, the time over which massive merging BBHs will emit 0.1-10 Hz GWs may be estimated as Peters (1964)
$T\approx 70\text{ s}\left(\frac{M}{10^{4}\text{ M}}_{\odot}\right)^{-\frac{5}{3}}\left(\frac{f_{GW}}{0.01\text{ Hz}}\right)^{-\frac{8}{3}}\frac{(1+q)^{2}}{q}.$ (3)
$10^{3}–10^{4}$ ${}_{\odot}$ will spend $100-10^{4}$ s in the APTA band before merging, while stellar-mass BBHs can remain for days to even years depending on the mass, allowing for long-term signal accumulation from these sources.
More unique astrophysical GW sources may also manifest in this frequency regime. Low frequency sources such as Type Ia Seitenzahl et al. (2015) and core-collapse supernovae Ott (2009), parabolic encounters Kocsis et al. (2006); Codazzo et al. (2023), white dwarf-neutron star mergers (Rebassa-Mansergas et al., 2019; Korol et al., 2022), and even some extreme-mass-ratio inspirals Finn and Thorne (2000b); Hughes (2000); Babak et al. (2017) should emit GWs in this band. With longer observations, APTA can tap lower frequencies. As such, APTA could be able to observe verification binaries to be used in LISA calibration Kupfer et al. (2018); Marsh (2011); Korol et al. (2017); Breivik et al. (2018); Tauris (2018), as well as probe the major astrophysics goals of LISA Amaro-Seoane et al. (2023).
Observations in the 0.1–10 Hz frequency band could also help constrain models of cosmic inflation through the detection of primordial gravitational waves Corbin and Cornish (2006); Smith et al. (2006); Kuroyanagi et al. (2009); Guzzetti et al. (2016). It is advantageous to explore this frequency range as opposed to LISA’s 0.11000.11000.1–1000.1 – 100 mHz Amaro-Seoane et al. (2017), since, in the latter, inflationary GWs may be overshadowed by binary white dwarf GWs, which disappear entirely above 0.25 Hz Ungarelli and Vecchio (2001); Farmer and Phinney (2003); Cutler and Holz (2009). Furthermore, heavy ($\lesssim 1$ $\mathrm{M}_{\odot}$) primordial-black-hole binaries, which could account for a small fraction of the dark matter in the form of massive compact halo objects (MACHOs), could be detected in the APTA band during their early inspiral Nakamura et al. (1997); Raidal et al. (2019). Another possibility of cosmological interest in the APTA band could be the direct measurement of the acceleration of the expansion of the universe and, therefore, the curvature parameter $\Omega_{k0}$ Nakamura and Chiba (1999); Seto et al. (2001).
III Pulsar Timing Array Sensitivity
[…]
When collecting data, GW detectors are subject to several sources of noise that complicate the identification of signals from astrophysical sources. The data stream received by the detector $\mathbb{s}(t)$ is the sum of a noise term $\mathbb{n}(t)$ and a (possible) signal term $\mathbb{h}(t)$:
$\mathbb{s}(t)=\mathbb{n}(t)+\mathbb{h}(t).$ (4)
The significance of the signal is determined by the signal-to-noise ratio (SNR). Taking a frequentist approach to GW signal processing, one defines a statistic $\mathfrak{s}$ by cross-correlating the timing residuals with a symmetric filter matrix $\mathbb{K}(t)$:
$\mathfrak{s}=\int\mathrm{d}t\int\mathrm{d}t^{\prime}\mathbb{s}^{T}(t)\mathbb{K}^{\dagger}(t-t^{\prime})\mathbb{s}(t^{\prime})=\int\mathrm{d}f\;\mathbb{\tilde{s}}(f)^{T}\mathbb{\tilde{K}}^{\dagger}(f)\mathbb{\tilde{s}}(f),$ (5)
where the tilde denotes a Fourier transform. The SNR $\rho$ is the ratio of the average of this statistic in the presence of a signal to its root mean square (rms) value with only noise:
$\rho^{2}=\frac{\langle\mathfrak{s}\rangle^{2}_{s=n+h}}{\langle\mathfrak{s}^{2}\rangle_{s=n}-\langle\mathfrak{s}\rangle_{s=n}^{2}}.$ (6)
In GW signal processing, it is common practice to look for the “matched” filter $\mathbb{\tilde{K}}(f)$ that maximizes $\rho$ in order to optimally extract the signal from the noise. $\mathbb{\tilde{K}}(f)$ and its corresponding $\rho$ may be evaluated by assuming Gaussian white noise and that the pulsars are far enough from each other to ensure the noise of individual pulsars is uncorrelated Moore et al. (2015a). In this case, for a monochromatic GW emitted by a slowly evolving compact object binary, the SNR is
$\rho^{2}=\frac{1}{2}N_{p}(N_{p}-1)\frac{8\chi^{4}h_{c}^{4}}{T}\int_{1/T}^{1/\delta t}\mathrm{d}f\;\frac{\sin^{4}{\left[\pi(f-f_{0})T\right]}}{\pi^{4}(f-f_{0})^{4}f^{4}S_{n}^{2}},$ (7)
where $N_{p}$ is the number of pulsars, $1/\delta t$ is the cadence of observations, $T$ is the total time of observation of the source, and $S_{n}=2\sigma^{2}\delta t$ is the noise power spectral density for white noise, where $\sigma$ is the rms error in the timing residuals. $\chi$ is a sky-averaged version of the redshift $z$ of pulsar signals. In general, $z$ is
$z=\frac{1}{2}\frac{\hat{p}^{i}\hat{p}^{j}}{1+\hat{\Omega}\cdot\hat{p}}\left(h_{ij}(t-|\mathbf{p}|(1+\hat{\Omega}\cdot\hat{p}),\hat{\Omega})-h_{ij}(t,\hat{\Omega})\right),$ (8)
where $t$ is the detection time, $\mathbf{p}$ is the position of the pulsar in an Earth-centered coordinate system, $\hat{\Omega}$ is the direction of propagation of the GW, and $h_{ij}$ is the GW perturbation to the background metric Anholm et al. (2009). In PTAs one often assumes that the pulsar term, i.e. the first term in Eq. 8, is zero Moore et al. (2015a). We discuss the implications of this in Section IV.2. Besides the sky-averaging, we average $z$ over GW polarization and source inclination angle so as not to overestimate the sensitivity of APTA, resulting in $\chi\approx 0.317$.
The integral in Eq. 7 can be analytically approximated in the high and low-frequency regimes Moore et al. (2015a). Choosing a detection-threshold value $\rho_{th}$ and inverting Eq. 7, the dimensionless characteristic GW strain $h_{c}$ in the high and low-frequency regimes is
$h_{c}^{\mathrm{HIGH}}(f)\approx\left(\frac{16\rho_{th}^{2}}{3\chi^{4}N_{p}(N_{p}-1)}\right)^{1/4}\sigma f\sqrt{\frac{\delta t}{T}}$ (9a)
$h_{c}^{\mathrm{LOW}}(f)\approx\frac{3\sqrt{\rho_{th}}}{2^{7/4}\chi\pi^{3}}\left(\frac{13}{N_{p}(N_{p}-1)}\right)^{1/4}\frac{\sigma}{T^{3}f^{2}}\sqrt{\frac{\delta t}{T}}\sec{\phi}$ (9b)
where $\phi$ is chosen to make $h_{c}^{\mathrm{HIGH}}(2/T)=h_{c}^{\mathrm{LOW}}(2/T)$. The total sensitivity is hence Moore et al. (2015a)
$h_{c}(f)\approx h_{c}^{\mathrm{HIGH}}(f)+h_{c}^{\mathrm{LOW}}(f).$ (10)
As this model shows, sensitivity curves fundamentally depend on the timing precision, the threshold SNR, the number of pulsars, and the cadence and total duration of observation.
IV Artificial Precision Timing Array
To construct a GW detector based on PTA principles, we envision distributing pulsing electromagnetic sources in the sky. These light sources need to have their periodic pulses timed to high precision so that GWs may be observed with them. This can be achieved with an array of atomic-clock-carrying satellites, which would play the same role as the astrophysical pulsars in regular PTAs. They could orbit around the Earth or the Sun. Alternatively, rather than Earth-based radio receivers, one could imagine a near-Earth receiving base to avoid atmospheric interference, as illustrated in Fig. 1. The details of this communication mechanism, including the electromagnetic pulse features and its use at detection, are beyond the scope of this paper, which focuses on clock requirements. However, this communication pathway must not degrade the required clock accuracy.
Figure 1: Cartoon illustration of the APTA concept. The detector is composed of an array of satellites communicating with a near-Earth central receiving station. The satellites, equipped with atomic clocks to allow for precision timing, play the role of the pulsars in this pulsar-timing-like GW detector. Objects are not drawn to scale, and the satellite array’s orbit may encircle some solar system planets, contrary to what the figure suggests.
The results of the model from Section III do not depend on the pulsar/satellite communication frequency because $f_{photon}\gg f_{GW}$ for reasonable signal frequency choices Isaacson (1968); Stappers et al. (2011). To observe 0.1100.1100.1–100.1 – 10 Hz GWs, radio waves would do, but visible light would give a pulse-to-GW frequency ratio comparable to that of regular PTAs. X-rays may also become a good choice given recent advances in the generation of high-frequency phase-stable pulses (Zhang et al., 2022, e.g.). We leave the optimal photon wavelength for APTA pulses as a subject for future investigation.
IV.1 Timing error and cadence
PTAs observing for a total time $T$ with a cadence of $1/\delta t$ are generally sensitive to GW frequencies in the range $(1/T)\lesssim f_{GW}\lesssim(1/\delta t)$ Moore et al. (2015b). Thus accessing the 0.1100.1100.1–100.1 – 10 Hz band can be achieved straightforwardly simply by searching for GW transients and continuous signals that last at least $\sim 10$ s and taking data at least every 0.1 s.
The lowest detectable characteristic strain depends linearly on the error in the PTA timing residuals $\sigma$ as indicated in Eqs. III. We estimate the attainable sensitivities for APTA given various timing errors consistent with present and future atomic clock precision. We defer the discussion of additional sources of error to Section V. The best atomic clocks at present can reach fractional frequency uncertainties $\xi=\Delta f_{clock}/f_{clock}\sim 10^{-18}$ in 1 s of averaging and $\sim 10^{-19}–10^{-20}$ in 1 h Bothwell et al. (2022); Oelker et al. (2019); Campbell et al. (2017). In the next few years, we may reach $\xi\sim 10^{-20}$ in 1 s averaging with current 3D optical lattice clock technology. In a decade, use of squeezing and increased atom number may allow $\xi\sim 10^{-21}$ Ludlow et al. (2015). Note that $\xi\propto t_{avg}^{-1/2}$, for $t_{avg}$ the time of averaging. Our averaging time is limited by cadence so $\xi$ over time $\delta t$ of averaging is
$\xi_{\delta t}=\xi_{1\text{ s}}\left(\frac{\delta t}{1\text{ s}}\right)^{-\frac{1}{2}},$ (11)
where $\xi_{1\text{ s}}$ is the fractional uncertainty in 1 s of averaging. We assume APTA’s clocks reach a stable optimal fractional uncertainty $\xi_{\delta t}$ after averging and use this value to compute the timing error as $\sigma=\xi_{\delta t}\delta t$.
APTA can be sensitive to much higher frequencies than regular PTAs because its precision timing can be carried out continuously, whereas PTAs are limited by radio telescope access. Continued observation can be achieved if the beacon satellites are locked to a central observing station. In this case, the cadence $1/\delta t$ is bounded by the frequency at which communication pulses can be stably emitted and received, thereby fundamentally limited by the atomic clocks’ frequency.
For communication purposes, an extremely efficient way to transfer the timing precision of optical atomic clocks to the radiofrequency domain is provided by optical frequency comb technology Udem et al. (2002); Diddams et al. (2020). It has the potential to generate precision APTA pulses at a carrier frequency of 10 MHz to 100 GHz. We also note that intense efforts are underway to miniaturize and ruggedize the technology of strontium optical lattice clocks Poli et al. (2014); Koller et al. (2017); Origlia et al. (2018); Takamoto et al. (2020). Progress has been fast, which makes the prospect of deploying precision optical lattice clocks within the next decade highly realistic.
In the simplified timing model described in Section III, lowering $\delta t$ always lowers the sensitivity. This is a feature of assuming white noise: when $\delta t\rightarrow 0$, the measurement becomes continuous, so noise does not remain at any frequency. The effects of more realistic noise will be considered in future work. The current state of the art for phase-stable X-ray pulses is $10^{-15}$ s Zhang et al. (2022) For this study, we use $\delta t=0.01$ s, an order of magnitude below the minimum requirement for accessing 10 Hz GWs. Therefore, to focus on clock uncertainty requirements, we utilize state-of-the-art low values of $\xi$ alongside conservative values of $\delta t$. The same sensitivities could be reached with higher $\xi$ by lowering $\delta t$. The determination of the optimal set of parameters for constructing APTA is left for subsequent studies.
IV.2 Distance to the satellite array
The resulting sensitivity curve in Eqs. III and 10 does not depend on the distance between the pulsars and the observer because the correlated timing residual at each pulsar is assumed to be zero across all pulsars. Recent studies have considered the sensitivity enhancement from modeling non-zero pulsar terms in the timing residual (see discussion in Ref. Zhu et al. (2016) and references therein), but we leave a more detailed treatment of the pulsar term for future work. This approximation assumes that pulsars are far enough from each other that their separation is much larger than the GW wavelength (for stochastic GW searches) and consequently the GW phase is not coherent across the beacons. While this may not be the case for APTA in practice, we use this model for the purposes of this proof of concept study.
In this initial study, we assume operators have perfect control of the satellites. Hence, increasing the number of satellites $N_{s}$ will always increase the detector’s sensitivity. For a realistic scenario, we may consider a handful of satellites, $N_{s}\approx 6$ in orbit at roughly 1-2 AU from the sun. Although the assumption of uncorrelated satellite timing residuals may start to break down at these distances, the detector can still work with non-zero correlations between the satellite timing residuals included in the real sensitivity model.
IV.3 APTA sensitivity curves
Fundamentally, the sensitivity model involves five free parameters: $\rho_{th}$, $\delta t$, $N_{s}$, $T$ and $\sigma$. Following Ref. Moore et al. (2015a), we choose $\rho_{th}=3$ as a standard value for the rest of our study. We use $\delta t=0.01$ s, which is an order of magnitude below the minimum requirement for accessing 10 Hz GWs. We fiducially choose $N_{s}=6$ as a reasonable number of satellites.
We assume that all the error in the timing residuals arises from the clock’s uncertainty, i.e. $\sigma=\xi\delta t$. Motivated by developments in clock technology (see Section IV.1), we choose three clock fractional uncertainties for our analysis: $\xi=10^{-17}$, the current record in 0.01 s of averaging Bothwell et al. (2022); Oelker et al. (2019); Campbell et al. (2017); $\xi=10^{-19}$, a realistic number for the next decade Ludlow et al. (2015); and $\xi=10^{-22}$, a high precision value included to establish an upper bound for $\xi$ in the foreseeable future.
In Fig. 2, we display APTA sensitivity curves alongside example GW strains of merging intermediate-mass BBHs and early inspiraling stellar-mass and primordial BBHs. With current atomic clock technology, represented by $\xi=10^{-17}$ Bothwell et al. (2022); Oelker et al. (2019); Campbell et al. (2017), APTA could observe $10^{3}$ $\mathrm{M}_{\odot}$ and $10^{4}$ $\mathrm{M}_{\odot}$ BBH mergers at Gpc distances, as well as briefly see $100$ $\mathrm{M}_{\odot}$ binary inspirals. These inspirals are more clearly observed by APTA with the $\xi=10^{-19}$ near-future clock technology. Progress to $\xi=10^{-22}$ would allow the detection of $0.1$ $\mathrm{M}_{\odot}$ primordial-black-hole binaries. Alternatively, the current clock state of the art could see these sources using a lower $\delta t$ or if they are closer.
Figure 2: Dimensionless characteristic GW strain as a function of frequency for different APTA clock precisions and binary black hole inspiral masses. The colored solid lines are APTA sensitivity curves for a total observation time $T=10$ s, pulse interval $\delta t=0.01$ s, number of satellites $N_{s}=6$, and threshold SNR $\rho_{th}=3$. The black lines represent binary black holes with mass ratio $q=1$ at a distance $D=1$ Gpc, with masses as indicated. The solid and dashed lines representing the binaries with total masses $M=10^{4}$ $\mathrm{M}_{\odot}$ and $M=10^{3}$ $\mathrm{M}_{\odot}$ end respectively at $f=1.3$ Hz and $f=13$ Hz, the approximate values for their mergers (see Eq. 2).
In Fig. 3, the sensitivity of APTA for the same three cases as in Fig. 2 is displayed together with the sensitivity curves of many current and future detectors, besides shaded areas corresponding to astrophysical GW sources Moore et al. ; Moore et al. (2015b). Sensitivity curve colors match those of Fig. 2, such that APTA-17 has $\xi=10^{-17}$, APTA-19 has $\xi=10^{-19}$, and APTA-22 has $\xi=10^{-22}$. With current clock technology Bothwell et al. (2022); Oelker et al. (2019); Campbell et al. (2017), APTA can surpass TianQin and LISA’s sensitivity in the dHz band and see some compact-binary inspirals. With near-future clock precision Ludlow et al. (2015), APTA may additionally access extreme-mass-ratio inspirals and GW bursts from Type Ia supernovae. Far future clocks with $\xi=10^{-22}$ can allow APTA to considerably outperform other proposed detectors DECIGO Kawamura et al. (2006) and BBO Phinney et al. (2004); Harry et al. (2006).
Figure 3: Dimensionless characteristic GW strain as a function of frequency for different astrophysical sources (shaded areas) and detectors (curves) Moore et al. ; Moore et al. (2015b). We display APTA for $[\delta t,T,\rho_{th},N_{s}]=[0.01\mathrm{s},10\mathrm{s},3,6]$. APTA-17 in purple has $\xi=10^{-17}$, APTA-19 in blue has $\xi=10^{-19}$, and APTA-22 in red has $\xi=10^{-22}$. The other detectors are Virgo Accadia et al. (2012), Advanced LIGO (aLIGO) Aasi et al. (2015), Einstein Telescope (ET) Punturo et al. (2010), Cosmic Explorer (CE) Reitze et al. (2019), TianQin Luo et al. (2016), the DECi-hertz Interferometer Gravitational wave Observatory (DECIGO) Kawamura et al. (2006), Big Bang Observer (BBO) Phinney et al. (2004); Harry et al. (2006), and LISA Amaro-Seoane et al. (2017).
V Timing model Considerations
Several additional effects must be taken into account in a realistic timing model. Although detailed calculation of these effects is beyond the scope of this study, we discuss some of them below.
Shapiro time delay, the change in the travel time of light between two points due to the presence of massive bodies Shapiro (1964), from various solar system objects will introduce delays of $2GM/c^{3}$, e.g. $\sim 10^{-5}$ for the Sun, $\sim 10^{-8}$ s for Jupiter, and $\sim 10^{-13}$ s for the Moon. Since the main asteroid belt has an estimated total mass of $18\times 10^{-10}$ $\mathrm{M}_{\odot}$ Krasinsky et al. (2002), asteroids may also contribute with delays $\sim 10^{-14}$ s. These known delays may need to be included in the timing model, depending on where the satellites are placed. The gravity gradient noise of asteroids, on the other hand, only contributes marginally at much lower frequencies ($10^{-9}–10^{-5}$ Hz) Fedderke et al. (2022); Sullivan et al. (2022), and so should not affect APTA.
After setting the satellites in orbit, their positions should be tracked with sufficient precision to account for their motion and added time delays. Additionally, it will be important to track disruptions to the network (e.g. a satellite being captured by a planet or colliding with an asteroid). Disruptions to one or a few elements will not necessarily prevent GW detection since the signals detected by the other satellites remain correlated. If further analyses reveal that this is an important concern, the detector can be designed with a redundantly large number of satellites.
Because of APTA’s small solar-system scale, some components of traditional PTA models, including parallax Deng and Finn (2011) as well as Einstein and Roemer delays (related to the use of a coordinate system centered in the solar system barycenter Hobbs et al. (2006)), are not as important. The solar wind Marsch (2006) contributes to the dispersion measure (DM), i.e. the integrated column density of free electrons between the observer and a light source. While accounting for the effects of the time-dependent DM on the ToAs is not straightforward Agazie et al. (2023b), this problem is easier for APTA than for PTAs, since the only contributions are from the solar system rather than the interstellar medium.
To first approximation, clock synchronization should not play an important role in the functioning of APTA. Regular PTAs are of course not synchronized clocks, and have already been proven successful by relying on the long-term timing precision of the individual pulsars. Still, in general, significant time precision losses may happen in the process of time transfer (Caldwell et al., 2023, e.g.). Preserving, transferring, and utilizing the satellite clocks’ precision is also mission-critical for success. Therefore, the generation of the electromagnetic pulses and their locking to the clock should preserve the accuracy of the time reference. This can be done with state-of-the-art frequency comb technology Udem et al. (2002); Diddams et al. (2020). Additionally, the pulses’ phase information and intensity must not degrade sufficiently during their travel to compromise the recovery of the timing information. The multi-beam detection scheme at the central location must also preserve the mission-critical relative timing information. Each of these interwoven problems will require its own individual technical analysis as the required clock accuracy only sets the performance specifications of the downstream systems engineering.
VI Conclusion
We have presented a proof-of-principle study into the use of an artificial solar-system-scale timing array to detect 0.1–10 Hz GWs, focusing on the required time reference precision. APTA would consist of $\sim 10$ satellites equipped with precision clocks orbiting and preferably communicating with a near-Earth base so the atmosphere is not in the path. The desired frequency band can be reached by targeting astrophysical signals at least 10 s long and ensuring that the beacons’ pulses are received at least every 0.1 s.
APTA’s sensitivity is fundamentally limited by $\sigma$, the error in the timing residuals. With the current state-of-the-art atomic clock technology Bothwell et al. (2022); Oelker et al. (2019); Campbell et al. (2017), APTA can achieve better dHz sensitivity than LISA and TianQin, allowing for detection of $10^{3}–10^{4}$ $\mathrm{M}_{\odot}$ black hole mergers and the early inspirals of some ground-based GW sources. With future clock technology available Ludlow et al. (2015), APTA could compete with and eventually surpass the sensitivity of other proposed detectors for the 0.1–10 Hz frequency range without the same challenges in high-precision interferometry they face. Then, it could detect most astrophysical sources in this band. These exciting prospects motivate the continued development of atomic clock technology and its application to GW detection.
While focusing on the accuracy of the APTA’s time reference, we have only discussed briefly the effects of additional noise sources, technical challenges, and technical developments required. This will require more detailed future work. Nevertheless, with conservative choices for measurement cadence $\delta t$, we have shown that high sensitivity can be still be achieved. Future studies may indicate that $\delta t$ can be lowered without straining APTA’s engineering requirements. With this and future papers laying the foundation, we propose APTA to observe uncharted territory of the gravitational-wave sky and gain new insight into high-energy astrophysics and cosmology.
Acknowledgements
The authors thank Columbia University in the City of New York and Stanford University for their generous support. The Columbia Experimental Gravity group is grateful for the generous support of Columbia University. LMBA is grateful for the Columbia Undergraduate Scholars Program Summer Enhancement Fellowship and the Columbia Center for Career Education Summer Funding Program. AS is grateful for the support of the Stanford University Physics Department Fellowship and the National Science Foundation Graduate Research Fellowship.